Bill's retirement plan consists of contributing $130 at the end of each month into an RRSP. If the plan earns 4.4% compounded quarterly for the first ten years and 6% compounded semiannually thereafter, how much money will be in Bill's RRSP at the end of a thirty year term?
Answer #1

$130 per month

1 year =130*12=1560

Quarterly compounding for first 10 years

A= I(1+r/m)^nm

1560(1+0.044/4)^10*4

1560(1+0.011)^40

2416.411

Semi annually

1560(1+0.06/2)20*2

1560(1+0.03)^40

5088.72